Alfalfa’s age triple age of his brother Ralfalfa. The sum of their ages is 40. Find the age of each.
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52776)

.

    x + 3x = 40


     4x    = 40


      x    = 40/4 = 10.


ANSWER.  The younger brother is 10 years old.  The older brother is 30 years old.

Solved.



Answer by Alan3354(69443)

A = 3R
A + R = 40